# Entropic Confinement and Mode Connectivity in Overparameterized Neural Networks

### v1

@@ Abstract
Modern neural networks exhibit a striking property: basins of attraction in the loss landscape are often connected by low-loss paths, yet optimization dynamics generally remain confined to a single convex basin (Baity-Jesi_2019; juneja2023linear) and rarely explore intermediate points. We resolve this paradox by identifying entropic barriers arising from the interplay between curvature variations along these paths and noise in optimization dynamics. Empirically, we find that curvature systematically rises away from minima, producing effective forces that bias noisy dynamics back toward the endpoints — even when the loss remains nearly flat. These barriers persist longer than energetic barriers, shaping the late-time localization of solutions in parameter space. Our results highlight the role of curvature-induced entropic forces in governing both connectivity and confinement in deep learning landscapes.
@@ 1 Introduction
Deep neural networks trained, in the overparametrized regime, exhibit a number of surprising and counterintuitive properties. One of the most striking is the observation that distinct solutions, found with standard optimization algorithms, are often connected by low-loss paths in parameter space (mode-con; pmlr-v80-draxler18a; frankle-LMC). Such mode connectivity results imply that the landscape is far less rugged than once assumed: minima that appear isolated are, in fact, linked by paths of low, nearly constant loss. At the same time, however, optimization dynamics display a seemingly contradictory behavior. Standard training with stochastic gradient descent (SGD), with or without momentum, converges to a well-defined minimum and rarely explores regions of parameter space corresponding to these paths (Baity-Jesi_2019).
@@ 1.1 Related Work
This behavior is attributed to the higher noise levels in gradient estimates, which act as a form of implicit regularization preventing convergence to sharp minima and instead favoring wider, flatter basins that generalize better (keskar2017on). Several works have formalized this intuition: jastrzębski2018finding and smith2018bayesian interpret the minibatch noise as an effective temperature that enables exploration and escape from sharp valleys; wei2019noiseaffectshessianspectrum and xie2021a further show that the resulting dynamics can be modeled as a stochastic process biased toward flat regions. Collectively, these studies demonstrate that curvature and generalization are deeply intertwined with the stochastic geometry of the optimization trajectory. We leverage these insights to conduct a deeper analysis of mode connectivity in neural network training. mode-con and pmlr-v80-draxler18a showed the existence of nonlinear paths of low-loss between different minima found by training with different random seeds. Subsequently, frankle-LMC showed that when the training dynamics of two networks are tied together early in training, the resulting minima, found after training is complete, are linearly connected by paths of low-loss. Follow-up work has analyzed in more depth how linear mode connectivity emerges (singh2024landscaping; zhou2023goinglinearmodeconnectivity), and such work has important implications for model merging (ainsworth2023git; singh2020fusion) and weight-space ensembling (izmailov2018swa; wortsman2021robust; gagnon-audet2023awe; pmlr-v162-wortsman22a).
@@ 2 Background: Entropic Forces and Curvature
In the analogy to deep learning, $V$ plays the role of the task loss function, while the noise $\xi$ arises from SGD noise due to minibatching and finite learning rate. To leading order one can identify $T\propto\eta/B$, for a learning rate $\eta$ and batch size $B$, (mandt2017stochastic; smith2020generalization; liu2021noise), though the precise relationship depends on local curvature and other details of the loss landscape (ziyin2021strength). In this simple motivating example we assume $\xi$ is white and Gaussian, even though in deep networks the SGD noise is neither perfectly white nor perfectly Gaussian. In this analogy, $y$ corresponds to soft modes (directions with nearly flat curvature where the loss hardly changes), and $x$ represents the stiff modes (directions associated with large eigenvalues of the Hessian).
@@
The resulting stochastic dynamics converge to a Boltzmann-like stationary distribution (gardiner2004handbook),
@@ 3 Methods
We begin by training a collection of image-classification models on CIFAR-10 using both Wide ResNet and ResNet architectures, each initialized from different random seeds to obtain a diverse set of distinct minima. For every pair of minima under study, we construct a low-loss path connecting them using the AutoNEB algorithm of pmlr-v80-draxler18a.
Once the MEP is obtained, we analyze its geometric and dynamical properties. First, we measure curvature along the path using several complementary Hessian-based statistics, allowing us to characterize how flatness varies between the endpoints and the interior of the path. Second, we study optimization dynamics constrained to the MEP by projecting SGD updates onto the nearest path segment. This controlled setting isolates how stochasticity interacts with curvature: models initialized along the MEP exhibit systematic drift toward flatter regions, revealing the action of entropic forces whose strength grows with the effective noise level. We then proceed to analyze this phenomenon at a finer scale by studying linearly connected minima, following the approach of frankle-LMC.
@@ 3.1 Training Details
Unless otherwise specified, all experiments are conducted on Wide ResNet-16-4 (zagoruyko2016wide) trained on the CIFAR-10 dataset (Krizhevsky09learningmultiple). Following standard practice (zagoruyko2016wide), we use stochastic gradient descent (SGD) with momentum $\beta=0.9$, weight decay regularization $w=5\times 10^{-4}$, and an initial learning rate of $\eta=0.1$. Models are trained for 200 epochs with a batch size of 256, and the learning rate is reduced by a factor of 5 at 30%, 60%, 80%, and 90% of the total training epochs. We apply mild data augmentation consisting of random horizontal flips and random crops with 4-pixel padding followed by resizing to the original $32\times 32$ resolution.
@@ 3.2 Minimum energy paths
To explore the structure of the loss landscape between different solutions, we identify low-loss connecting paths using the Automatic Nudged Elastic Band (AutoNEB) algorithm introduced by pmlr-v80-draxler18a. In brief, the algorithm initializes $k$ intermediate pivots along the straight-line path between two minima and optimizes their positions such that they evolve as if connected by elastic springs, while minimizing the loss orthogonal to the path.
Since the loss along the straight segments between pivots may still be high, AutoNEB dynamically adds new pivots whenever the loss along a segment exceeds a predefined threshold. This adaptive refinement ensures a smooth, low-loss path is found. Following pmlr-v80-draxler18a, we refer to such paths as minimum energy paths (MEPs), by analogy with physical systems. Importantly, the optimization dynamics is such that it does not change the length of the segments composing the MEP, so when a new pivot is inserted between two existing pivots, the resulting segments remain shorter than the original. In the following, the relative position along the MEP is reported in terms of pivot index, normalized by the total number of pivots. Note that this parameterization does not reflect the actual metric distance along the path, as the pivot density and length are non-uniform, see Figure 7 in Appendix A.3.
@@ 3.3 Curvature Measures
We estimate the maximum eigenvalue of the Hessian, $\lambda_{\max}(\mathcal{H})$, using the power iteration method (see, e.g., yao2020pyhessian). Crucially, this method requires only Hessian–vector products, which can be computed efficiently via automatic differentiation in $\mathcal{O}(N)$ time. The update rule for the power method is:
@@ 3.3.1 A Note on Reparameterization
dinh2017sharp showed that symmetries in the architecture of networks allow deep networks to be re-parameterized without changing the function computed by the network. While this observation potentially makes the Hessian a poor tool for studying generalization, we note that when considering SGD optimization dynamics it is still the Hessian and not a reparameterization invariant measure that governs the dynamics of the system. Particularly, for any symmetry $T_{\alpha}$ that leaves the function computed by the network the same, we have that $\gradient_{\alpha}L(T_{\alpha}\theta)=0$, and so flat directions induced by a symmetry do not induce a gradient.
@@ 4.1 Entropic Confinement
One might argue that the increase in sharpness along the MEP is simply a consequence of the decreasing loss or an effective reduction in learning rate, especially given prior work suggesting a relationship between sharpness and learning rate (cohen2021gradient). We argue that this is not the case: while the loss drops between the first and second pivots, it then remains approximately constant along the rest of the MEP. In contrast, both sharpness metrics—maximum eigenvalue and trace of the Hessian—continue to rise. This indicates that the observed increase in curvature is not merely a byproduct of lower loss or implicit regularization, but rather reflects a genuine change in the geometry of the optimization landscape.
@@ 4.2 Linear mode connectivity
Figure 5: Entropic barriers are relevant later in training. (A) Linear mode connectivity schematic (frankle-LMC). We train a network to epoch $k$, then produce two new networks via different data ordering, and measure the loss along a linear path. (B) The average loss along such a path goes down as $k$ increases, decreasing rapidly with $k$. (C) Top: The loss profile along linear paths for various $k$. Bottom: The curvature profile, measured by the maximum Hessian eigenvalue, for various $k$. (D) We plot the instability (The relative change along the path) of the loss and the curvature. For small $k$, the loss exhibits larger instability, while for larger $k$, the curvature exhibits larger instability.
Although we have argued that entropic forces separate the low-loss region of parameter space into regions effectively confined by entropic barriers, we have not yet addressed how and when these confined regions are chosen along the course of training. In this section we will take steps towards answering this question through the lens of linear mode connectivity. Following the methods of frankle-LMC, we train $M$ networks with a shared data order up until epoch $k$, which we will call the splitting epoch. After epoch $k$, each of the $M$ networks sees an independent ordering of the data and can then potentially move away from its “siblings,” the other M1M-1 networks. The sibling networks are then trained until convergence. All networks trained in this section use the ResNet-20 architecture (he2015deepresiduallearningimage), unless otherwise noted.
The crucial observation in frankle-LMC is that once $k$ becomes sufficiently large, the sibling networks become connected by linear paths of low loss, implying that they converge to the same region of parameter space. Interestingly, $k$ does not have to be very large compared to the number of epochs required for convergence before linear mode connectivity is observed. In Figure 5, we reproduce these experiments, and measure the curvature along the linear, low-loss paths between converged siblings. Similarly to the nonlinear case (Section 4.1), we see a bump in the curvature along these paths. Additionally, we notice that these entropic barriers persist for larger values of $k$ than their energetic counterparts, implying that entropic forces contribute relatively more to the final stages of the model’s localization to a region of parameter space. To see this, we plot the instability as a function of $k$ (Figure 5D). The instability measures the relative change, max/min, in the metric (loss or curvature) along the linear path. For small values of $k$, the loss has the larger instability, while for larger values of $k$ the curvature exhibits greater instability.
@@ 5 Discussion
Our results provide new insight into the global geometry of the loss landscape. While prior work has emphasized that minima are often connected by low-loss paths, forming a single broad “valley” of solutions (mode-con; frankle-LMC), our findings reveal that these paths are not flat when entropic forces are taken into consideration (Figure 2). Instead, they exhibit systematic increases in curvature away from their endpoints, producing localized “bumps” in sharpness. This observation refines the valley picture: the basin of low-loss solutions is structured by curvature variations that give rise to entropic barriers.
@@
Our analysis of linear mode connectivity further shows that entropic forces play an important role late in training. As the splitting epoch increases, energetic barriers along linear paths decrease, but curvature barriers persist for longer into training (Figure 5, Figure 6). This suggests a two-phase picture of training: early dynamics are dominated by energetic forces that drive the model into a low-loss basin, while later on entropic forces become more relevant. Our experiments have important implications for late-time dynamics of deep network training, basin selection, and parameter-space ensembling techniques. In particular, earlier work including altintas2025butterfly, demonstrates that the final basin a network ends up in is highly sensitive to perturbations in the weights, especially early in training. These findings suggest that even small contributions from entropic forces could have an outsized effect on the model’s long-term fate.
@@
We posit that generalizing minima may be effectively disconnected from overfit minima via entropic barriers. Entropic barriers could make paths to such regions effectively inaccessible: even when overfitting solutions are connected to flatter ones by low-loss paths, entropic forces could shield the generalizing solutions by repelling SGD away from regions of parameter space that do not generalize. Our results suggest that this is a promising avenue for future work. In fact, there is evidence that models in similar convex basins of attraction share generalization properties (juneja2023linear).
@@
Our work also provides a new lens through which to view weight-space ensembling techniques. The study of global loss landscape features, such as mode connectivity (pmlr-v80-draxler18a; frankle-LMC), has been crucial in developing methods like Stochastic Weight Averaging (SWA) (izmailov2018swa; wortsman2021robust). Our findings suggest a more nuanced picture of the global landscape: techniques like SWA may be averaging minima that, while energetically connected within a single low-loss valley, may be effectively disconnected by the entropic barriers we observe. This would imply that the SWA solution cannot be easily found by diffusive optimization dynamics at the bottom of a valley in the loss landscape. A valuable avenue for future work would be to analyze the connectivity properties of these averaged minima to better understand how weight-space averaging is able to construct solutions with favorable generalization properties.

### v2

@@ Abstract
Modern neural networks exhibit a striking property: basins of attraction in the loss landscape are often connected by low-loss paths, yet optimization dynamics generally remain confined to a single convex basin (Baity-Jesi et al., 2019; Juneja et al., 2023) and rarely explore intermediate points. We resolve this paradox by identifying entropic barriers arising from the interplay between curvature variations along these paths and noise in optimization dynamics. Empirically, we find that curvature systematically rises away from minima, producing effective forces that bias noisy dynamics back toward the endpoints — even when the loss remains nearly flat. These barriers persist longer than energetic barriers, shaping the late-time localization of solutions in parameter space. Our results highlight the role of curvature-induced entropic forces in governing both connectivity and confinement in deep learning landscapes.
@@ 1 Introduction
Deep neural networks trained, in the overparametrized regime, exhibit a number of surprising and counterintuitive properties. One of the most striking is the observation that distinct solutions, found with standard optimization algorithms, are often connected by low-loss paths in parameter space (Garipov et al., 2018; Draxler et al., 2018; Frankle et al., 2020). Such mode connectivity results imply that the landscape is far less rugged than once assumed: minima that appear isolated are, in fact, linked by paths of low, nearly constant loss. At the same time, however, optimization dynamics display a seemingly contradictory behavior. Standard training with stochastic gradient descent (SGD), with or without momentum, converges to a well-defined minimum and rarely explores regions of parameter space corresponding to these paths (Baity-Jesi et al., 2019).
@@ 1.1 Related Work
This behavior is attributed to the higher noise levels in gradient estimates, which act as a form of implicit regularization preventing convergence to sharp minima and instead favoring wider, flatter basins that generalize better (Keskar et al., 2017). Several works have formalized this intuition: jastrzębski2018finding and Smith and Le (2018) interpret the minibatch noise as an effective temperature that enables exploration and escape from sharp valleys; Wei and Schwab (2019) and Xie et al. (2021) further show that the resulting dynamics can be modeled as a stochastic process biased toward flat regions. Collectively, these studies demonstrate that curvature and generalization are deeply intertwined with the stochastic geometry of the optimization trajectory. We leverage these insights to conduct a deeper analysis of mode connectivity in neural network training. Garipov et al. (2018) and Draxler et al. (2018) showed the existence of nonlinear paths of low-loss between different minima found by training with different random seeds. Subsequently, Frankle et al. (2020) showed that when the training dynamics of two networks are tied together early in training, the resulting minima, found after training is complete, are linearly connected by paths of low-loss. Follow-up work has analyzed in more depth how linear mode connectivity emerges (Singh et al., 2024; Zhou et al., 2023), and such work has important implications for model merging (Ainsworth et al., 2023; Singh and Jaggi, 2020) and weight-space ensembling (Izmailov et al., 2018; Wortsman et al., 2021; Gagnon-Audet et al., 2023; Wortsman et al., 2022).
@@ 2 Background: Entropic Forces and Curvature
In the analogy to deep learning, $V$ plays the role of the task loss function, while the noise $\xi$ arises from SGD noise due to minibatching and finite learning rate. To leading order one can identify $T\propto\eta/B$, for a learning rate $\eta$ and batch size $B$, (Mandt et al., 2017; Smith et al., 2020; Liu et al., 2021), though the precise relationship depends on local curvature and other details of the loss landscape (Ziyin et al., 2021). In this simple motivating example we assume $\xi$ is white and Gaussian, even though in deep networks the SGD noise is neither perfectly white nor perfectly Gaussian. In this analogy, $y$ corresponds to soft modes (directions with nearly flat curvature where the loss hardly changes), and $x$ represents the stiff modes (directions associated with large eigenvalues of the Hessian).
@@
The resulting stochastic dynamics converge to a Boltzmann-like stationary distribution (Gardiner and others, 2004),
@@ 3 Methods
We begin by training a collection of image-classification models on CIFAR-10 using both Wide ResNet and ResNet architectures, each initialized from different random seeds to obtain a diverse set of distinct minima. For every pair of minima under study, we construct a low-loss path connecting them using the AutoNEB algorithm of Draxler et al. (2018).
Once the MEP is obtained, we analyze its geometric and dynamical properties. First, we measure curvature along the path using several complementary Hessian-based statistics, allowing us to characterize how flatness varies between the endpoints and the interior of the path. Second, we study optimization dynamics constrained to the MEP by projecting SGD updates onto the nearest path segment. This controlled setting isolates how stochasticity interacts with curvature: models initialized along the MEP exhibit systematic drift toward flatter regions, revealing the action of entropic forces whose strength grows with the effective noise level. We then proceed to analyze this phenomenon at a finer scale by studying linearly connected minima, following the approach of Frankle et al. (2020).
@@ 3.1 Training Details
Unless otherwise specified, all experiments are conducted on Wide ResNet-16-4 (Zagoruyko and Komodakis, 2016) trained on the CIFAR-10 dataset (Krizhevsky, 2009). Following standard practice (Zagoruyko and Komodakis, 2016), we use stochastic gradient descent (SGD) with momentum $\beta=0.9$, weight decay regularization $w=5\times 10^{-4}$, and an initial learning rate of $\eta=0.1$. Models are trained for 200 epochs with a batch size of 256, and the learning rate is reduced by a factor of 5 at 30%, 60%, 80%, and 90% of the total training epochs. We apply mild data augmentation consisting of random horizontal flips and random crops with 4-pixel padding followed by resizing to the original $32\times 32$ resolution.
@@ 3.2 Minimum energy paths
To explore the structure of the loss landscape between different solutions, we identify low-loss connecting paths using the Automatic Nudged Elastic Band (AutoNEB) algorithm introduced by Draxler et al. (2018). In brief, the algorithm initializes $k$ intermediate pivots along the straight-line path between two minima and optimizes their positions such that they evolve as if connected by elastic springs, while minimizing the loss orthogonal to the path.
Since the loss along the straight segments between pivots may still be high, AutoNEB dynamically adds new pivots whenever the loss along a segment exceeds a predefined threshold. This adaptive refinement ensures a smooth, low-loss path is found. Following Draxler et al. (2018), we refer to such paths as minimum energy paths (MEPs), by analogy with physical systems. Importantly, the optimization dynamics is such that it does not change the length of the segments composing the MEP, so when a new pivot is inserted between two existing pivots, the resulting segments remain shorter than the original. In the following, the relative position along the MEP is reported in terms of pivot index, normalized by the total number of pivots. Note that this parameterization does not reflect the actual metric distance along the path, as the pivot density and length are non-uniform, see Figure 7 in Appendix A.3.
@@ 3.3 Curvature Measures
We estimate the maximum eigenvalue of the Hessian, $\lambda_{\max}(\mathcal{H})$, using the power iteration method (see, e.g., Yao et al. (2020)). Crucially, this method requires only Hessian–vector products, which can be computed efficiently via automatic differentiation in $\mathcal{O}(N)$ time. The update rule for the power method is:
@@ 3.3.1 A Note on Reparameterization
Dinh et al. (2017) showed that symmetries in the architecture of networks allow deep networks to be re-parameterized without changing the function computed by the network. While this observation potentially makes the Hessian a poor tool for studying generalization, we note that when considering SGD optimization dynamics it is still the Hessian and not a reparameterization invariant measure that governs the dynamics of the system. Particularly, for any symmetry $T_{\alpha}$ that leaves the function computed by the network the same, we have that $\gradient_{\alpha}L(T_{\alpha}\theta)=0$, and so flat directions induced by a symmetry do not induce a gradient.
@@ 4.1 Entropic Confinement
One might argue that the increase in sharpness along the MEP is simply a consequence of the decreasing loss or an effective reduction in learning rate, especially given prior work suggesting a relationship between sharpness and learning rate (Cohen et al., 2021). We argue that this is not the case: while the loss drops between the first and second pivots, it then remains approximately constant along the rest of the MEP. In contrast, both sharpness metrics—maximum eigenvalue and trace of the Hessian—continue to rise. This indicates that the observed increase in curvature is not merely a byproduct of lower loss or implicit regularization, but rather reflects a genuine change in the geometry of the optimization landscape.
@@ 4.2 Linear mode connectivity
Figure 5: Entropic barriers are relevant later in training. (A) Linear mode connectivity schematic (Frankle et al., 2020). We train a network to epoch $k$, then produce two new networks via different data ordering, and measure the loss along a linear path. (B) The average loss along such a path goes down as $k$ increases, decreasing rapidly with $k$. (C) Top: The loss profile along linear paths for various $k$. Bottom: The curvature profile, measured by the maximum Hessian eigenvalue, for various $k$. (D) We plot the instability (The relative change along the path) of the loss and the curvature. For small $k$, the loss exhibits larger instability, while for larger $k$, the curvature exhibits larger instability.
Although we have argued that entropic forces separate the low-loss region of parameter space into regions effectively confined by entropic barriers, we have not yet addressed how and when these confined regions are chosen along the course of training. In this section we will take steps towards answering this question through the lens of linear mode connectivity. Following the methods of Frankle et al. (2020), we train $M$ networks with a shared data order up until epoch $k$, which we will call the splitting epoch. After epoch $k$, each of the $M$ networks sees an independent ordering of the data and can then potentially move away from its “siblings,” the other M1M-1 networks. The sibling networks are then trained until convergence. All networks trained in this section use the ResNet-20 architecture (He et al., 2015), unless otherwise noted.
The crucial observation in Frankle et al. (2020) is that once $k$ becomes sufficiently large, the sibling networks become connected by linear paths of low loss, implying that they converge to the same region of parameter space. Interestingly, $k$ does not have to be very large compared to the number of epochs required for convergence before linear mode connectivity is observed. In Figure 5, we reproduce these experiments, and measure the curvature along the linear, low-loss paths between converged siblings. Similarly to the nonlinear case (Section 4.1), we see a bump in the curvature along these paths. Additionally, we notice that these entropic barriers persist for larger values of $k$ than their energetic counterparts, implying that entropic forces contribute relatively more to the final stages of the model’s localization to a region of parameter space. To see this, we plot the instability as a function of $k$ (Figure 5D). The instability measures the relative change, max/min, in the metric (loss or curvature) along the linear path. For small values of $k$, the loss has the larger instability, while for larger values of $k$ the curvature exhibits greater instability.
@@ 5 Discussion
Our results provide new insight into the global geometry of the loss landscape. While prior work has emphasized that minima are often connected by low-loss paths, forming a single broad “valley” of solutions (Garipov et al., 2018; Frankle et al., 2020), our findings reveal that these paths are not flat when entropic forces are taken into consideration (Figure 2). Instead, they exhibit systematic increases in curvature away from their endpoints, producing localized “bumps” in sharpness. This observation refines the valley picture: the basin of low-loss solutions is structured by curvature variations that give rise to entropic barriers.
@@
Our analysis of linear mode connectivity further shows that entropic forces play an important role late in training. As the splitting epoch increases, energetic barriers along linear paths decrease, but curvature barriers persist for longer into training (Figure 5, Figure 6). This suggests a two-phase picture of training: early dynamics are dominated by energetic forces that drive the model into a low-loss basin, while later on entropic forces become more relevant. Our experiments have important implications for late-time dynamics of deep network training, basin selection, and parameter-space ensembling techniques. In particular, earlier work including Altıntaş et al. (2025), demonstrates that the final basin a network ends up in is highly sensitive to perturbations in the weights, especially early in training. These findings suggest that even small contributions from entropic forces could have an outsized effect on the model’s long-term fate.
@@
We posit that generalizing minima may be effectively disconnected from overfit minima via entropic barriers. Entropic barriers could make paths to such regions effectively inaccessible: even when overfitting solutions are connected to flatter ones by low-loss paths, entropic forces could shield the generalizing solutions by repelling SGD away from regions of parameter space that do not generalize. Our results suggest that this is a promising avenue for future work. In fact, there is evidence that models in similar convex basins of attraction share generalization properties (Juneja et al., 2023).
@@
Our work also provides a new lens through which to view weight-space ensembling techniques. The study of global loss landscape features, such as mode connectivity (Draxler et al., 2018; Frankle et al., 2020), has been crucial in developing methods like Stochastic Weight Averaging (SWA) (Izmailov et al., 2018; Wortsman et al., 2021). Our findings suggest a more nuanced picture of the global landscape: techniques like SWA may be averaging minima that, while energetically connected within a single low-loss valley, may be effectively disconnected by the entropic barriers we observe. This would imply that the SWA solution cannot be easily found by diffusive optimization dynamics at the bottom of a valley in the loss landscape. A valuable avenue for future work would be to analyze the connectivity properties of these averaged minima to better understand how weight-space averaging is able to construct solutions with favorable generalization properties.
@@ 6 Conclusion
Ethics Statement.
We do not foresee any direct ethical concerns arising from this work. Our study focuses on better understanding optimization in machine learning, without direct deployment in sensitive application domains. We note that we have used language models to polish the text of this manuscript in places.
Reproducibility Statement.
We have provided descriptions of all algorithms, models, and experimental setups in the main text and appendix. Training procedures and dataset details are documented to facilitate replication. When the author list is unblinded, we will release our codebase to enable full reproducibility of our results.